# BiSSL: Enhancing the Alignment Between Self-Supervised Pretraining and Downstream Fine-Tuning via Bilevel Optimization

### v4

@@ 3.4 Training Algorithm and Pipeline
1:Input: Backbone and pretext head initializations ${\bf\it\theta}$, ${\bf\it\phi}_{P}$, ${\bf\it\phi}_{D}$. Training objectives $\mathcal{L}^{P}$, $\mathcal{L}^{D}$. Optimizers $\text{opt}_{P}$, $\text{opt}_{D}$. Regularization Weight $\lambda\in\mathbb{R}_{+}$. Number of training stage alternations $T\in\mathbb{N}$ with upper and lower-level iterations $N_{U},N_{L}\in\mathbb{N}$.
@@
7:         Compute $\mathbf{g}_{{\bf\it\phi}_{P}}=\nabla_{{\bf\it\phi}}\mathcal{L}^{P}({\bf\it\theta}_{P},{\bf\it\phi})|_{{\bf\it\phi}={\bf\it\phi}_{P}}$
@@
14:         Approximate $\mathbf{v}_{\text{IJ}}\approx\quantity[\frac{1}{\lambda}\nabla_{{\bf\it\theta}}^{2}\mathcal{L}^{P}\quantity({\bf\it\theta},{\bf\it\phi}_{P})|_{{\bf\it\theta}={\bf\it\theta}_{P}}+I_{L}]^{-1}\mathbf{v}$. $\triangleright$ Use CG
@@
Algorithm 1 outlines the proposed training algorithm, which iteratively alternates between solving the lower-level (3) and upper-level (2) optimization problems in BiSSL. The lower-level training optimizes the pretext task objective, while additionally including the gradient of the regularization term $r$ for the backbone parameter updates, complying with (3). For the upper-level training, the gradient with respect to the backbone parameters as represented by the left term on the right-hand side in (7), is approximated using the CG method. We remark that allowing $N_{U}>1$ deviates from conventional BLO implementations. However, as documented in Section C.3 of Appendix C, this modification was found to be beneficial for downstream performance. The algorithm is applicable to any common pretext and downstream tasks, with the only supplementary requirement being that the pretext task objective is twice-differentiable.
Figure 1 illustrates the proposed training pipeline with BiSSL alongside the conventional SSL pipeline. Pretext pretraining on the unlabeled dataset $\mathcal{D}^{P}$ provides initializations of ${\bf\it\theta}$ and ${\bf\it\phi}_{P}$, after which the downstream head ${\bf\it\phi}_{D}$ is fitted on top of the frozen backbone ${\bf\it\theta}$ using the downstream dataset $\mathcal{D}^{D}$. BiSSL training is then conducted as outlined in Algorithm 1, yielding an updated backbone ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})$, which serves as the initialization for final supervised fine-tuning on the downstream task.
As detailed in Section A.2 in Appendix A, ${\bf\it\theta}^{\ast}_{P}({\bf\it\theta}_{D})$ must satisfy the stationary condition $\nabla_{{\bf\it\theta}}\quantity(\mathcal{L}^{P}({\bf\it\theta},{\bf\it\phi}_{P})+\lambda r({\bf\it\theta}_{D},{\bf\it\theta}))|_{{\bf\it\theta}={\bf\it\theta}^{\ast}_{P}({\bf\it\theta}_{D})}=\mathbf{0}$ to justify the explicit expression of the IJ in (6). This theoretical requirement motivates the placement of BiSSL after the pretraining stage, as a converged initialization ${\bf\it\theta}$ achieved from pretraining ensures the stationary condition is roughly satisfied at the outset. A similar argument applies to ${\bf\it\phi}_{D}$, since a randomly initialized downstream head typically causes rapid updates to ${\bf\it\theta}_{D}$ during early fine-tuning, likely violating the stationary assumption due to its coupling with ${\bf\it\theta}_{P}$ through the regularization objective $r$.
@@ 4.1 Datasets
The ImageNet-1K [36] dataset devoid of labels is used for self-supervised pretraining throughout the main experiments. For downstream fine-tuning and evaluation, we leverage a varied set of natural image classification datasets that encompass a wide array of tasks, including general image classification, fine-grained recognition across species and objects, scene understanding, and texture categorization. The datasets include Food 101 [52], CIFAR10 [53], CIFAR100 [53], Caltech-UCSD Birds-200-2011 (CUB200) [54], SUN397 scene dataset [55], StanfordCars [56], FGVC Aircraft [57], PASCAL VOC 2007 [58], Describable Textures Dataset (DTD) [59], Oxford-IIIT Pets [60], Caltech-101 [61] and Oxford 102 Flowers [62]. All downstream datasets are split into training, validation, and test partitions, with details on how these assignments are made in Section B.1 of Appendix B.
@@ 4.2.1 Baseline Setup
We conduct experiments using two different types of pretext tasks: SimCLR [1] and Bootstrap Your Own Latent (BYOL) [4]. For SimCLR we use a temperature of $0.5$ and a ResNet-50 [35] backbone. On top of the backbone, a projection head is applied, consisting of two fully connected layers with batch normalization [63] and ReLU [64] followed by a single linear layer. Each layer consists of $256$ neurons. For BYOL, we use constant target decay rate of $0.9995$, with the backbone and projection head architectures identical to those of SimCLR. The additional BYOL-specific prediction head uses an architecture identical to the projection head. The remaining details in this paragraph apply to both SimCLR and BYOL.
[…]
@@
[…]
A random grid search of $100$ hyper-parameter configurations for the base learning rates and weight decays is conducted, where one model is fine-tuned for each configuration. Base learning rates and weight decays are log-uniformly sampled over the ranges of $10^{-4}$ to $1$ and $10^{-5}$ to $10^{-2}$, respectively. Validation data accuracy is evaluated after each epoch. The hyper-parameter configuration yielding the best balance between high validation accuracy and low validation loss is considered the optimal hyper-parameter configuration. [^2] In certain scenarios during the experiments, the configuration that achieved the highest validation accuracy also yielded a notably higher relative validation loss. To ensure better generalizability, an alternative configuration with a more favorable trade-off was selected in these cases. The corresponding optimal hyper-parameters for each downstream dataset are documented in Table 3 of Appendix B. For subsequent evaluation on the test data, we train $10$ models with different random seeds, each using the considered optimal hyper-parameter configurations. During the training of each respective model, the model parameters are stored after each epoch if the top-1 validation accuracy (or 11-point mAP for the VOC07 dataset) has increased compared to the previous highest top-1 validation accuracy achieved during training. Top-1 and top-5 test data accuracies (or 11-point mAP for the VOC07 dataset) are evaluated for each of the $10$ models, from which the calculated means and standard deviations of these accuracies are documented.
@@ 4.2.2 BiSSL Setup
In this section, we detail each stage of the proposed training pipeline for BiSSL, as outlined in the right part of Figure 1.
Pretext Task
@@
The training setup for the downstream head warm-up closely mirrors the fine-tuning setup of Section 4.2.1. The main difference is that only the linear downstream head is fitted on top of the now frozen backbone obtained from the pretext warm-up. Learning rates and weight decays are initially selected based on those listed in Table 3, with adjustments made as needed when preliminary testing indicated a potential for improved convergence. These values are provided in Table 4 in Appendix B. The authors recognize that more optimal hyper-parameter configurations may exist and leave further exploration of this for future refinement. The downstream head warm-up is conducted for $20$ epochs with a constant learning rate.
@@
The training configuration for the lower-level primarily follows the setup described for pretext pretraining in Section 4.2.1, with the modifications outlined here. As specified in (3), the lower-level loss function is the sum of the pretext task objective $\mathcal{L}^{P}$ (e.g. the NT-Xent loss for SimCLR [1]) and the regularization term $r(\boldsymbol{\theta}_{D},\boldsymbol{\theta}_{P})=\frac{1}{2}||{\bf\it\theta}_{D}-{\bf\it\theta}_{P}||_{2}^{2}$, aligning with Algorithm 1. Based on early experiments, the regularization weight $\lambda=0.001$ was selected, as it appeared to strike a well-balanced compromise between the convergence rates of both the lower- and upper-level objectives. The lower-level is trained for the equivalent of approximately $8$ conventional pretraining epochs, with further details provided in the composite configuration paragraph. Accordingly the linear learning rate scheduler warm-up is adjusted to range over $10\cdot N_{L}$ training steps, with $N_{L}$ being the number of lower-level iterations as specified in Algorithm 1.
@@
The upper-level training stage largely mirrors the downstream training setup described in Section 4.2.1, and again, only the differences are addressed here. The weight decays and base learning rates are set to match those obtained from the downstream head warm-up detailed in Table 4 of Appendix B. To approximate the upper-level gradient in  (7), the conjugate gradient method [47, 48] is employed, with implementation details covered in Section B.3 of Appendix B.
@@
As outlined in Algorithm 1, both lower- and upper-level backbone parameters ${\bf\it\theta}_{P}$ and ${\bf\it\theta}_{D}$ are initialized with the backbone parameters obtained during the pretext warm-up, and the training procedure alternates between solving the lower- and upper-level optimization problems. In this experimental setup, the lower-level performs $N_{L}=20$ gradient steps before alternating to the upper-level, which then conducts $N_{U}=8$ gradient steps. A total of $T=500$ training stage alternations are executed. With the ImageNet dataset and the current batch size of $1024$, there are a total of $1251$ training batches without replacement. Consequently, the $T=500$ training stage alternations correspond to roughly $8$ conventional pretext epochs, a negligible additional training load compared to the $500$ pretext epochs used for the full pretraining process. Additionally, gradient normalization is employed on gradients exceeding $\ell_{2}$-norms of $10$.
@@
Table 1: Comparison of classification accuracies between the conventional SSL pipeline and the proposed BiSSL pipeline. Accuracies significantly higher from their counterparts are marked in bold.
@@
Only FT $75.0\pm 0.1$ $96.1\pm 0.2$ $79.5\pm 0.2$ $49.6\pm 0.3$ $49.3\pm 0.2$ $78.0\pm 0.3$ $52.0\pm 1.0$ $71.0\pm 0.1$ $60.3\pm 0.9$ $73.2\pm 0.3$ $85.7\pm 0.5$ $82.6\pm 0.3$
BiSSL+FT $\mathbf{75.6\pm 0.1}$ $96.4\pm 0.1$ $\mathbf{80.0\pm 0.1}$ $\mathbf{59.2\pm 0.2}$ $\mathbf{51.4\pm 0.2}$ $78.0\pm 0.3$ $\mathbf{55.3\pm 0.5}$ $\mathbf{71.8\pm 0.1}$ $\mathbf{63.4\pm 0.3}$ $\mathbf{79.6\pm 0.3}$ $\mathbf{88.1\pm 0.3}$ $\mathbf{85.6\pm 0.2}$
[…]
@@
Only FT $75.3\pm 0.2$ $96.4\pm 0.1$ $80.1\pm 0.2$ $52.7\pm 0.4$ $47.9\pm 0.3$ $76.9\pm 0.2$ $51.6\pm 0.7$ $69.3\pm 0.1$ $59.5\pm 0.4$ $77.9\pm 0.3$ $86.6\pm 0.4$ $82.2\pm 0.5$
BiSSL+FT $75.4\pm 0.1$ $96.4\pm 0.1$ $\mathbf{80.6\pm 0.1}$ $\mathbf{60.1\pm 0.1}$ $\mathbf{50.2\pm 0.2}$ $\mathbf{78.1\pm 0.3}$ $\mathbf{57.1\pm 0.4}$ $\mathbf{71.1\pm 0.1}$ $\mathbf{62.2\pm 0.1}$ $\mathbf{81.9\pm 0.1}$ $\mathbf{88.3\pm 0.3}$ $\mathbf{86.5\pm 0.1}$
[…]
4.3 Downstream Task Performance
We first assess the impact of using BiSSL compared to the conventional self-supervised training pipeline by evaluating classification accuracies on the various specified downstream datasets. Table 1 report means and standard deviations of top-1 classification accuracies (or the 11-point mAP on the VOC2007 dataset), comparing results obtained from the conventional SSL pipeline with those achieved using the BiSSL pipeline. Table 7 of Section C.2 in Appendix C outlines the corresponding top-5 accuracies. The results reveal BiSSL significantly improves downstream performance on $10$ out of $12$ datasets, with no single result showing a decline in performance compared to the baseline. Moreover, classification accuracies achieved via BiSSL exhibit greater stability, as evidenced by consistently lower or comparable variances relative to the baseline.
Table 2: $\text{AP}_{50}$-scores for object detection on the VOC07+12 dataset using a Faster-RCNN with a R50-C4 backbone.
SimCLR BYOL
Only FT $59.6\pm 0.3$ $54.6\pm 0.5$
BiSSL+FT $\mathbf{60.6\pm 0.1}$ $\mathbf{56.7\pm 0.4}$
Object Detection

### v5

@@ 3.4 Training Algorithm and Pipeline
1:Input: Backbone and head initializations $\bf\it\theta$}, ${\bf\it\phi}_{P}$, ${\bf\it\phi}_{D}$. Training objectives $\mathcal{L}^{P}$, $\mathcal{L}^{D}$. Optimizers $\text{opt}_{P}$, $\text{opt}_{D}$. Regularization Weight $\lambda\in\mathbb{R}_{+}$. Number of training stage alternations $T\in\mathbb{N}$ with upper and lower-level iterations $N_{U},N_{L}\in\mathbb{N}$.
@@
7:   Compute $\mathbf{g}_{{\bf\it\phi}_{P}}=\nabla_{{\bf\it\phi}}\mathcal{L}^{P}({\bf\it\theta}_{P},{\bf\it\phi})|_{{\bf\it\phi}={\bf\it\phi}_{P}}$ .
@@
14:   Approximate $\mathbf{v}_{\text{IJ}}\approx\quantity[\frac{1}{\lambda}\nabla_{{\bf\it\theta}}^{2}\mathcal{L}^{P}\quantity({\bf\it\theta},{\bf\it\phi}_{P})|_{{\bf\it\theta}={\bf\it\theta}_{P}}+I_{L}]^{-1}\mathbf{v}$. $\triangleright$ Use CG to approximate
@@
Our proposed training algorithm iteratively alternates between solving the lower-level (3) and upper-level (2) optimization problems in BiSSL. The lower-level training optimizes the pretext task objective, while additionally including the gradient of the regularization term $r$, complying with (3). The upper-level stage computes gradients with respect to the backbone, corresponding to the left-hand term in (7), which are approximated using the CG method. Algorithm 1 details the proposed training procedure. We remark that allowing $N_{U}>1$ deviates from conventional BLO training setups. However, as we document in Section C.3, this modification proved more efficient and beneficial for downstream performance. The algorithm is applicable to any common pretext and downstream tasks.
Figure 1 illustrates the proposed training pipeline with BiSSL alongside the conventional SSL pipeline. Pretext pretraining on the unlabeled dataset $\mathcal{D}^{P}$ provides initializations of $\bf\it\theta$} and ${\bf\it\phi}_{P}$, after which the downstream head ${\bf\it\phi}_{D}$ is fitted on top of the frozen backbone $\bf\it\theta$} using the downstream dataset $\mathcal{D}^{D}$. BiSSL training is then conducted as outlined in Algorithm 1, yielding an updated backbone ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})$, which serves as the initialization for final supervised fine-tuning on the downstream task. We emphasize that the objective $\mathcal{L}^{D}$ in the upper-level of BiSSL is identical to the downstream fine-tuning objective, and similarly that the objective $\mathcal{L}^{P}$ in the lower-level is identical to the pretext pretraining objective.
Section A.1 notes that ${\bf\it\theta}^{\ast}_{P}({\bf\it\theta}_{D})$ must satisfy the stationary condition $\nabla_{{\bf\it\theta}}\quantity(\mathcal{L}^{P}({\bf\it\theta},{\bf\it\phi}_{P})+\lambda r({\bf\it\theta},{\bf\it\theta}_{D}))|_{{\bf\it\theta}={\bf\it\theta}^{\ast}_{P}({\bf\it\theta}_{D})}=\mathbf{0}$ to justify the explicit expression of the IJ in (6). This requirement naturally aligns with using a backbone that has already converged from self-supervised pretraining, as such an initialization roughly satisfies the stationary condition at the outset and also makes the smoothness assumption on $\mathcal{L}^{P}$, mentioned in Section A.1, more reasonable. A similar consideration applies to ${\bf\it\phi}_{D}$, since a randomly initialized downstream head can lead to rapid early updates to ${\bf\it\theta}_{D}$ during fine-tuning, potentially violating the stationary assumption due to its coupling with ${\bf\it\theta}_{P}$ through the regularization objective $r$.
@@ 4.1 Datasets
The ImageNet-1K (Deng et al., 2009) dataset, devoid of labels, is used for self-supervised pretraining throughout the main experiments. For downstream fine-tuning and evaluation, we leverage a varied set of natural image classification datasets that encompass a wide array of tasks, including general image classification, fine-grained recognition across species and objects, scene understanding, and texture categorization. The datasets include Food 101 (Bossard et al., 2014), CIFAR10 (Krizhevsky, 2012), CIFAR100 (Krizhevsky, 2012), Caltech-UCSD Birds-200-2011 (CUB200) (Wah et al., 2011), SUN397 scene dataset (Xiao et al., 2010), StanfordCars (Yang et al., 2015), FGVC Aircraft (Maji et al., 2013), PASCAL VOC 2007 (Everingham et al., 2010), Describable Textures Dataset (DTD) (Cimpoi et al., 2014), Oxford-IIIT Pets (Parkhi et al., 2012), Caltech-101 (Li et al., 2022a) and Oxford 102 Flowers (Nilsback and Zisserman, 2008). Additionally, we use the combined VOC07+12 (Everingham et al., 2010) dataset for both object detection and semantic segmentation, and the fine-grained partition of Cityscapes (Cordts et al., 2016) for semantic segmentation. All downstream datasets are split into training, validation, and test partitions, with details on how these assignments are made in Section B.1.
@@ 4.2.1 Baseline Setup
We first outline each stage of the baseline setup following the standard SSL pipeline on the left of Figure 1.
@@
We conduct experiments using two different types of pretext tasks: SimCLR (Chen et al., 2020) and Bootstrap Your Own Latent (BYOL) (Grill et al., 2020). For SimCLR we use a temperature of $0.5$ and a ResNet-50 (He et al., 2016) backbone. On top of the backbone, a projection head is applied, consisting of two fully connected layers with batch normalization (Ioffe and Szegedy, 2015) and ReLU (Agarap, 2018) followed by a single linear layer. Each layer consists of $256$ neurons. For BYOL, we use constant target decay rate of $0.9995$, with the backbone and projection head architectures identical to those of SimCLR. The additional BYOL-specific prediction head uses an architecture identical to the projection head. The remaining details in this paragraph apply to both SimCLR and BYOL.
The image augmentation scheme follows the approach used in Bardes et al. (2022), with minor modifications: The image size is set to $96\crossproduct 96$ instead of $224\crossproduct 224$, and the minimal ratio of the random crop is adjusted accordingly to $0.5$ instead of $0.08$. The implementation of the LARS optimizer (You et al., 2017) from Bardes et al. (2022) is employed, with a “trust” coefficient of $0.001$, a weight decay of $10^{-6}$ and a momentum of $0.9$. The learning rate increases linearly during the first $10$ epochs, reaching a peak base learning rate of $4.8$, followed by a cosine decay with no restarts (Loshchilov and Hutter, 2017) for the remaining epochs. A batch size of $1024$ is used and, unless otherwise specified, pretraining is conducted for $500$ epochs.
@@
For downstream fine-tuning, a single linear layer is attached to the output of the pretrained backbone. The training procedure utilizes the cross-entropy loss, the SGD optimizer with a momentum of $0.9$, and a cosine decaying learning rate scheduler without restarts (Loshchilov and Hutter, 2017). Fine-tuning is conducted for $400$ epochs with a batch size of $256$. An augmentation scheme similar to the fine-tuning augmentation scheme in Bardes et al. (2022) is employed during training, where images are center cropped and resized to $96\crossproduct 96$ pixels with a minimal crop ratio of $0.5$, followed by random horizontal flips. Learning rates and weight decays for each downstream datasets are selected via a random hyper-parameter grid search. After hyperparameters are determined, we train $10$ models with different random seeds, and report the mean and standard deviation of the top-1 test accuracies. Details on the hyperparameter search and evaluation protocol are provided in Section B.2.
@@ 4.2.2 BiSSL Setup
This section detail each stage of the training pipeline involving BiSSL, as illustrated on the right of Figure 1.
Pretext Task Training
@@
The training setup for the downstream head warm-up closely mirrors the fine-tuning setup of Section 4.2.1. The main difference is that only the linear downstream head is fitted on top of the now frozen backbone obtained from the pretext warm-up. Learning rates and weight decays are initially selected based on those listed in Table 5, with adjustments made as needed when preliminary testing indicated a potential for improved convergence. These values are provided in Table 6 of Appendix B. The authors recognize that more optimal hyper-parameter values may exist and leave further exploration of this for future refinement. The downstream head warm-up is run for $20$ epochs with a constant learning rate.
@@
[…]
@@
The upper-level training stage largely mirrors the downstream training setup described in Section 4.2.1, and we again only list the differences. The weight decays and base learning rates are set to match those obtained from the downstream head warm-up detailed in Table 6 of Appendix B. To approximate the upper-level gradient in  (7), the conjugate gradient method (Nazareth, 2009; Shewchuk, 1994) is employed, with implementation details covered in Section B.3.
Table 1: Comparison of classification accuracies obtained through fine-tuning without and with preparatory BiSSL training. Statistically significant improvements over the counterpart are highlighted in bold.
SimCLR:
FT $75.0\pm 0.1$ $96.1\pm 0.2$ $79.5\pm 0.2$ $49.6\pm 0.3$ $49.3\pm 0.2$ $77.9\pm 0.3$ $52.0\pm 1.0$ $71.0\pm 0.1$ $60.4\pm 0.9$ $73.5\pm 0.7$ $85.8\pm 0.6$ $82.6\pm 0.3$
BiSSL+FT $\mathbf{76.2\pm 0.1}$ $\mathbf{96.3\pm 0.1}$ $\mathbf{79.8\pm 0.2}$ $\mathbf{55.7\pm 0.2}$ $\mathbf{50.9\pm 0.1}$ $77.9\pm 0.3$ $\mathbf{55.9\pm 0.3}$ $\mathbf{71.5\pm 0.1}$ $\mathbf{64.2\pm 0.4}$ $\mathbf{78.3\pm 0.3}$ $\mathbf{87.2\pm 0.3}$ $\mathbf{84.1\pm 0.2}$
Avg Diff $\mathbf{+1.2}$ $\mathbf{+0.2}$ $\mathbf{+0.4}$ $\mathbf{+6.1}$ $\mathbf{+1.6}$ $0.0$ $\mathbf{+3.9}$ $\mathbf{+0.5}$ $\mathbf{+3.8}$ $\mathbf{+4.8}$ $\mathbf{+1.4}$ $\mathbf{+1.5}$
BYOL:
FT $75.3\pm 0.3$ $96.4\pm 0.1$ $80.1\pm 0.2$ $52.7\pm 0.4$ $47.9\pm 0.3$ $76.9\pm 0.2$ $51.6\pm 0.7$ $69.3\pm 0.1$ $59.5\pm 0.4$ $77.9\pm 0.3$ $86.6\pm 0.4$ $82.1\pm 0.5$
BiSSL+FT $\mathbf{76.2\pm 0.1}$ $96.4\pm 0.1$ $\mathbf{80.4\pm 0.1}$ $\mathbf{57.1\pm 0.3}$ $\mathbf{49.6\pm 0.1}$ $\mathbf{77.5\pm 0.2}$ $\mathbf{57.2\pm 0.5}$ $\mathbf{70.8\pm 0.1}$ $\mathbf{61.7\pm 0.2}$ $\mathbf{80.1\pm 0.3}$ $\mathbf{87.2\pm 0.2}$ $\mathbf{85.4\pm 0.2}$
Avg Diff $\mathbf{+0.9}$ $0.0$ $\mathbf{+0.3}$ $\mathbf{+4.4}$ $\mathbf{+1.7}$ $\mathbf{+0.6}$ $\mathbf{+5.6}$ $\mathbf{+1.5}$ $\mathbf{+2.2}$ $\mathbf{+2.2}$ $\mathbf{+0.6}$ $\mathbf{+3.3}$
@@
As outlined in Algorithm 1, both lower- and upper-level backbone parameters ${\bf\it\theta}_{P}$ and ${\bf\it\theta}_{D}$ are initialized with the backbone parameters obtained during the pretext warm-up, and the training procedure alternates between solving the lower- and upper-level optimization problems. In this experimental setup, the lower-level performs $N_{L}=20$ gradient steps before alternating to the upper-level, which then conducts $N_{U}=8$ gradient steps. A total of $T=500$ training stage alternations are executed. With the ImageNet dataset and the current batch size of $1024$, there are a total of $1251$ training batches without replacement. Consequently, the $T=500$ training stage alternations correspond to roughly $8$ conventional pretext epochs, a negligible additional training load compared to the $500$ pretext epochs used for the full pretraining process. Section B.4 outlines further details on how data batches are handled during training. Lastly, gradient normalization is employed on gradients exceeding $\ell_{2}$-norms of $10$.
@@
4.3 Downstream Task Performance
Table 1 reports means and standard deviations of top-1 classification accuracies (or 11-point mAP on the VOC07 dataset), comparing results obtained from direct fine-tuning the pretrained backbone with those achieved by applying BiSSL prior to fine-tuning. We mark statistically significantly different accuracies with bold-font, with details on how the statistical significance tests are executed in Section B.6. Table 11 of Section C.2 outlines the corresponding top-5 accuracies. For both pretext tasks, BiSSL significantly improves downstream performance on $11$ out of $12$ datasets, with no single result showing a decline in performance. Moreover, accuracies from BiSSL exhibit lower or similar variances than the baseline. Smaller gains on tasks like CIFAR10, VOC07, and Caltech101 may reflect their coarse-grained nature and input distributions being more similar to the pretext task, whereas fine-grained tasks with more distinct input distributions that require more specialized representations like CUB200, Aircrafts, and Pets, exhibit greater gains from BiSSL.
Table 2: Comparison of $AP_{50}$ scores for object detection and mIoU scores for semantic segmentation obtained by fine-tuning without and with preparatory BiSSL training. Scores statistically significantly higher from their counterparts are marked in bold.
Detection ($AP_{50}$) Semantic Seg. (mIoU)
VOC07+12 VOC07+12 CityScapes
@@ 4.3 Downstream Task Performance
FT $53.1\pm 0.3$ $72.7\pm 0.6$ $47.8\pm 0.4$
BiSSL+FT $\mathbf{54.4\pm 0.4}$ $\mathbf{74.4\pm 0.6}$ $\mathbf{50.7\pm 0.4}$
Avg Diff $\mathbf{+1.3}$ $\mathbf{+1.7}$ $\mathbf{+2.9}$
@@
FT $52.6\pm 0.7$ $71.7\pm 0.6$ $48.6\pm 0.5$
BiSSL+FT $\mathbf{53.5\pm 0.3}$ $\mathbf{73.2\pm 0.4}$ $\mathbf{49.7\pm 0.4}$
Avg Diff $\mathbf{+0.9}$ $\mathbf{+1.5}$ $\mathbf{+1.1}$
4.3.1 Object Detection and Semantic Segmentation